# Multi-Agent Dynamic Relational Reasoning for Social Robot Navigation

### v1

@@ VI-D2 Social Robot Navigation
To simulate the behaviors of the robot and pedestrians, we employ holonomic kinematics. The action space for each agent, including the robot, at time is represented by their desired velocities along the and axes, denoted as . This action space is continuous, with the robot’s maximum speed capped at . We operate under the assumption that all agents can instantaneously achieve their desired velocities and maintain these velocities for a duration of seconds. The position of an agent at any time is updated based on the velocities according to the update rule, which ensures realistic motion patterns for the robot and human agents.
@@ VI-E1 Quantitative Analysis
ETH-UCY and SDD datasets. We adopted the standard setting for training and testing our model to predict trajectories in the future 4.8 seconds (12 time steps) based on the historical observations of 3.2 seconds (8 time steps). Our approach reduces minADE / minFDE by 14.3% / 16.7% and 20.0% / 22.7% on the ETH-UCY (overall) and ETH-UCY (difficult) datasets compared to the strongest baseline, respectively. Particularly, our method leads to a larger improvement for the difficult subset, indicating its advantage in modeling complex social interactions in dense scenarios. Our approach also achieves state-of-the-art performance on the SDD dataset.
NBA dataset. We predicted the trajectories in the future 4.0 seconds (10 time steps) based on the historical observations of 2.0 seconds (5 time steps). Our approach reduces minADE / minFDE by 20.5% / 25.1% on the NBA dataset compared to the strongest baseline.
@@ VI-E2 Ablative Analysis
Number |
2 |
5 |
11 |
20 |
30 |
40 |
110 |
minADE |
1.22 0.01 |
1.16 0.02 |
1.28 0.04 |
1.29 0.03 |
1.31 0.04 |
1.32 0.05 |
1.34 0.07 |
Number |
1 |
2 |
3 |
4 |
5 |
6 |
7 |
minADE |
1.27 0.01 |
1.23 0.02 |
1.16 0.02 |
1.19 0.02 |
1.23 0.04 |
1.27 0.05 |
1.29 0.04 |
@@
Group-wise relational reasoning. We show the effectiveness of group-wise relational reasoning by comparing the SCG, SHG, and SCG+SHG models. In all the experiments, SHG performs worse than SCG, which implies that group-wise relational reasoning through hypergraphs alone cannot sufficiently model the interactions between agents. This is because pairwise interactions still dominate the multi-agent dynamics of the agents in either the same or different groups, which cannot be entirely ignored. Meanwhile, effective pairwise reasoning may help identify group behavior patterns. Among these model variants, SCG+SHG achieves the best performance, which implies the effectiveness of integrating pairwise and group-wise relational reasoning. SCG+SHG reduces minADE by 4.2%, 6.7%, 5.8%, and 7.8% on the ETH-UCY (overall), ETH-UCY (difficult), SDD, and NBA datasets, respectively. As shown in Fig. 6(a) and 6(b), SCG+SHG also leads to a significant improvement in crowd simulation and the gap becomes larger as the prediction horizon increases.
@@
Dynamic Relational Reasoning. We show the effectiveness of dynamic relational reasoning by comparing the SCG+SHG and DCG+DHG models. DCG+DHG performs better in all the experiments. In the crowd simulation where the group evolves significantly, the dynamic relational reasoning mechanism enhances model performance. Since SCG+SHG cannot predict the evolution of hyperedges, the model must infer an overall group topology for the whole prediction, which limits the flexibility of reasoning. DCG+DHG reduces minADE by 8.7%, 10.7%, 20.8%, and 12.2% on ETH-UCY (overall), ETH-UCY (difficult), SDD, and NBA datasets, respectively.
@@
Smoothness regularization. We show the effectiveness of smoothness regularization by comparing the DCG+DHG and DCG+DHG+SM models. Fig. 6(a) and 6(b) show that DCG+DHG+SM achieves consistently better and more stable performance. The reason is that the smoothness regularization encourages smoother evolution of graphs and hypergraphs over time and restrains abrupt changes, which leads to a more stable training process and faster convergence. It also encourages the model to learn the temporal dependency between consecutively evolving relations. Applying the smoothness regularization loss during training reduces minADE by 4.8%, 8.0%, 7.4%, and 6.6% on the ETH-UCY (overall), ETH-UCY (difficult), SDD, and NBA datasets, respectively.
@@
Sharpness regularization. We show the effectiveness of the sharpness regularization by comparing DCG+DHG+SM and DCG+DHG+SM+SH models. Fig. 6(a) illustrates that the sharpness regularization further enhances the prediction performance in crowd simulation. Fig. 6(b) shows that DCG+DHG+SM+SH has a smaller variance in minADE indicating that sharpness regularization also improves the model robustness to random initializations. Applying the sharpness regularization during training reduces minADE by 5.0%, 4.3%, 4.5%, and 3.9% on the ETH-UCY (overall), ETH-UCY (difficult), SDD, and NBA datasets, respectively.
@@
Sparsity regularization. DCG+DHG+SM+SH+SP has the smallest prediction error among all the model settings consistently. DCG+DHG+SM+SH+SP tends to infer sparser hypergraphs which learn more distinguishable group memberships for each agent. The model is encouraged to learn to recognize sparse pairwise relations and clear group assignments, which creates a strong inductive bias in the message passing process through the sparse relational structures. Applying sparsity regularization during training reduces minADE by 5.3%, 9.1%, 3.6%, and 5.7% on the ETH-UCY (overall), ETH-UCY (difficult), SDD, and NBA datasets, respectively.
@@
Hyperedge numbers. In Table IV, we compare the model performance with different hyperedge (i.e., group) numbers on the NBA dataset. It shows that setting 5 as the maximum number of hyperedges achieves the best performance based on a hyperparameter search. The model may not capture sufficient group-wise relations with only 2 hyperedges. At the same time, too many hyperedges (e.g., 11 and 110) may lead to overly dense group-wise relations that cannot capture clear group behaviors. In Table I-III, the reported numbers of SHG show the best performance after doing a hyperparameter search. Based on the experimental results, we find that using a very large number of hyperedges in SHG leads to serious overfitting. Moreover, the standard deviation of the prediction error increases as the maximum hyperedge number increases because a larger number of hyperedges may lead to more local optima, which enlarges the randomness of the model performance. These observations imply that although pairwise relations are a special type of group-wise relations where only two agents are involved, it is not sufficient to just use a large number of hyperedges to cover all the relations, which motivates our design of the integrated pairwise and group-wise relational reasoning.
@@
Edge and hyperedge types. In Table V, we compare the model performance with different numbers of edge and hyperedge types (i.e., relation types) on the NBA dataset. It shows that setting 3 as the edge and hyperedge type number achieves the best performance based on a hyperparameter search. The model may not capture a sufficient variety of pairwise or group-wise relations with only 2 edge and hyperedge types while too many types (e.g., 6) may lead to overfitting. We can see in Table V that the standard deviation of the prediction error increases as the type number increases because a larger variety of relations may lead to more local optima, which enlarges the randomness of the model performance.
@@
Attention mechanisms. To validate the effectiveness of the attention mechanisms in Eq. (5) and (10), we replace the attention weights in DCG+DHG+SM+SH+SP with equal weights, which shows that adopting the attention mechanism reduces minADE by 14.3%, 16.7%, 13.8%, and 11.5% on the ETH-UCY (overall), ETH-UCY (difficult), SDD, and NBA datasets, respectively. This implies the benefits of modeling different contributions of agents to group-wise relations and the influence of different groups on a certain agent.
@@ VI-E3 Qualitative Analysis
Methods |
SR | CR | TR | NT | PL | ITR | SD | GI |
ORCA [66] |
82.311.82 |
17.181.71 |
0.510.11 |
14.910.09 |
24.710.88 |
12.110.32 |
0.350.02 |
21.351.31 |
RL (no prediction) |
85.212.35 |
14.582.14 |
0.210.08 |
16.350.16 |
21.130.20 |
6.940.27 |
0.340.03 |
9.870.20 |
RL (constant velocity) |
89.621.57 |
10.261.41 |
0.120.05 |
15.850.06 |
20.990.11 |
3.350.11 |
0.370.02 |
8.710.12 |
CrowdNav++ [13] |
93.371.33 |
6.541.24 |
0.090.03 |
14.780.08 |
19.580.13 |
3.150.12 |
0.420.02 |
8.510.14 |
Ours (w/o DHG+GR) |
95.391.41 |
4.531.33 |
0.080.03 |
14.820.08 |
19.570.14 |
3.110.11 |
0.410.01 |
8.660.18 |
Ours (w/o GR) |
97.210.58 |
2.710.55 |
0.080.03 |
14.720.10 |
19.630.15 |
2.980.10 |
0.420.01 |
8.110.12 |
Ours |
97.820.52 | 2.110.50 | 0.070.03 |
14.880.11 |
19.820.13 |
2.910.10 | 0.470.01 | 3.250.08 |
Oracle |
98.030.41 |
1.900.35 |
0.070.02 |
14.120.05 |
19.410.08 |
2.850.08 |
0.480.01 |
3.170.04 |
@@ VI-F1 Quantitative Analysis
By employing our trajectory prediction model with relational reasoning, the social navigation method significantly outperforms the rule-based method ORCA [66], the non-predictive method RL (no prediction), and predictive baselines RL (constant velocity) and CrowdNav++ [13] across all the evaluation metrics. The improvement is evident in achieving a higher success rate (SR), lower collision rate (CR) and timeout rate (TR), and shorter navigation time (NT) and path lengths (PL). Meanwhile, our method notably enlarges the social distance from humans and reduces improper intrusions into human personal or group spaces, as indicated by SD, ITR, and GI. This enables the robot to find safer navigation trajectories and minimize interference with pedestrians. Moreover, our method achieves comparable performance to the Oracle method which has access to the ground truth human trajectories, showing the effectiveness of our prediction model.
@@ VI-F2 Ablative Analysis
Group-wise reasoning. Ours (w/o DHG+GR) in Table VI represents a variant of our method with only pairwise reasoning and hence it disables group-based reward. Due to the lack of understanding of group relations, this variant leads to a higher CR and lower SD because the prediction model may not predict group behaviors accurately, which degrades the safety of robot navigation. Meanwhile, it causes a higher ITR and GI during navigation, which may make the surrounding humans less comfortable. Although it achieves a slightly shorter NT and PL (i.e., higher efficiency), the improvement is brought by the cost of social appropriateness, which is not desired.
Group reward. Ours (w/o GR) in Table VI represents a variant of our method without applying the group-based reward for RL policy learning. It still uses group-wise reasoning for trajectory prediction. The results show that without explicitly considering group relations in the reward function design, the robot policy still leads to a higher CR, ITR, GI, and a lower SD, although it achieves a comparable SR to our complete method. The NT and PL of our method are slightly larger because the policy tends to choose a safer and socially compliant trajectory which may take detours during navigation, which is acceptable and even desired in practice.
@@ VII Conclusion
In this paper, we present a systematic dynamic relational reasoning framework that infers pairwise and group-wise relations, which proves to be effective for multi-agent trajectory prediction and social robot navigation. In addition to inferring the conventional graphs that model pairwise relations between interacting agents, we propose to infer relation hypergraphs to model group-wise relations that exist widely in real-world scenarios such as human crowds. Different hyperedge types represent distinct group-wise relations. We employ a dynamic reasoning mechanism to capture the evolution of underlying relational structures. We propose effective mechanisms to regularize the smoothness of relation evolution, the sharpness of the learned relations, and the sparsity of the learned graphs and hypergraphs, which not only enhances training stability but also reduces prediction error. The proposed approach infers reasonable relations and achieves state-of-the-art performance on both synthetic crowd simulations and multiple benchmark datasets. Moreover, we present a deep reinforcement learning framework for social robot navigation that incorporates relational reasoning and trajectory prediction in a systematic way. It achieves state-of-the-art performance in terms of safety, efficiency, and social appropriateness compared with advanced baselines. A limitation of this work is that although our method outperforms the strongest baseline by a large margin, it still cannot achieve a zero collision rate in very dense scenarios. Therefore, in future work, we will investigate how to enable a provably safety guarantee. We will also conduct hardware experiments to further validate our approach.

### v2

@@ VI-D2 Social Robot Navigation
To simulate the behaviors of the robot and pedestrians, we employ holonomic kinematics. The action space for each agent, including the robot, at time is represented by their desired velocities along the and axes, denoted as . This action space is continuous, with the robot’s maximum speed at . Humans are controlled by ORCA [67] and interact with surrounding humans and obstacles. We operate under the assumption that all agents can instantaneously achieve their desired velocities and maintain these velocities for a duration of seconds. The position of an agent at any time is updated based on the velocities according to the update rule, which ensures realistic motion patterns for the robot and human agents.
@@ VI-E1 Quantitative Analysis
ETH-UCY and SDD datasets
We adopted the standard setting for training and testing our model to predict trajectories in the future 4.8 seconds (12 time steps) based on the historical observations of 3.2 seconds (8 time steps). Our approach reduces minADE20 / minFDE20 by 14.3% / 16.7% and 20.0% / 22.7% on the ETH-UCY (overall) and ETH-UCY (difficult) datasets compared to the strongest baseline, respectively. Particularly, our method leads to a larger improvement for the difficult subset, indicating its advantage in modeling complex social interactions in dense scenarios. Our approach also achieves state-of-the-art performance on the SDD dataset.
NBA dataset
We predicted the trajectories in the future 4.0 seconds (10 time steps) based on the historical observations of 2.0 seconds (5 time steps). Our approach reduces minADE20 / minFDE20 by 20.5% / 25.1% on the NBA dataset compared to the strongest baseline.
@@ VI-E2 Ablative Analysis
Group-wise relational reasoning
@@
We show the effectiveness of group-wise relational reasoning by comparing the SCG, SHG, and SCG+SHG models. In all the experiments, SHG performs worse than SCG, which implies that group-wise relational reasoning through hypergraphs alone cannot sufficiently model the interactions between agents. This is because pairwise interactions still dominate the multi-agent dynamics of the agents in either the same or different groups, which cannot be entirely ignored. Meanwhile, effective pairwise reasoning may help identify group behavior patterns. Among these model variants, SCG+SHG achieves the best performance, which implies the effectiveness of integrating pairwise and group-wise relational reasoning. SCG+SHG reduces minADE20 by 4.2%, 6.7%, 5.8%, and 7.8% on the ETH-UCY (overall), ETH-UCY (difficult), SDD, and NBA datasets, respectively. As shown in Fig. 6(a) and 6(b), SCG+SHG also leads to a significant improvement in crowd simulation and the gap becomes larger as the prediction horizon increases.
Dynamic Relational Reasoning
@@
We show the effectiveness of dynamic relational reasoning by comparing the SCG+SHG and DCG+DHG models. DCG+DHG performs better in all the experiments. In the crowd simulation where the group evolves significantly, the dynamic relational reasoning mechanism enhances model performance. Since SCG+SHG cannot predict the evolution of hyperedges, the model must infer an overall group topology for the whole prediction, which limits the flexibility of reasoning. DCG+DHG reduces minADE20 by 8.7%, 10.7%, 20.8%, and 12.2% on ETH-UCY (overall), ETH-UCY (difficult), SDD, and NBA datasets.
Number | 2 | 5 | 11 | 20 | 30 | 40 | 110 |
minADE20 | 1.22 0.01 | 1.16 0.02 | 1.28 0.04 | 1.29 0.03 | 1.31 0.04 | 1.32 0.05 | 1.34 0.07 |
Number | 1 | 2 | 3 | 4 | 5 | 6 | 7 |
minADE20 | 1.27 0.01 | 1.23 0.02 | 1.16 0.02 | 1.19 0.02 | 1.23 0.04 | 1.27 0.05 | 1.29 0.04 |
Smoothness regularization
@@
We show the effectiveness of smoothness regularization by comparing the DCG+DHG and DCG+DHG+SM models. Fig. 6(a) and 6(b) show that DCG+DHG+SM achieves consistently better and more stable performance. The reason is that the smoothness regularization encourages smoother evolution of graphs and hypergraphs over time and restrains abrupt changes, which leads to a more stable training process and faster convergence. It also encourages the model to learn the temporal dependency between consecutively evolving relations. Applying the smoothness regularization loss during training reduces minADE20 by 4.8%, 8.0%, 7.4%, and 6.6% on the ETH-UCY (overall), ETH-UCY (difficult), SDD, and NBA datasets, respectively.
Sharpness regularization
@@
We show the effectiveness of the sharpness regularization by comparing DCG+DHG+SM and DCG+DHG+SM+SH models. Fig. 6(a) illustrates that the sharpness regularization further enhances the prediction performance in crowd simulation. Fig. 6(b) shows that DCG+DHG+SM+SH has a smaller variance in minADE20 indicating that sharpness regularization also improves the model robustness to random initializations. Applying the sharpness regularization during training reduces minADE20 by 5.0%, 4.3%, 4.5%, and 3.9% on the ETH-UCY (overall), ETH-UCY (difficult), SDD, and NBA datasets, respectively.
Sparsity regularization
@@
DCG+DHG+SM+SH+SP has the smallest prediction error among all the model settings consistently. DCG+DHG+SM+SH+SP tends to infer sparser hypergraphs which learn more distinguishable group memberships for each agent. The model is encouraged to learn to recognize sparse pairwise relations and clear group assignments, which creates a strong inductive bias in the message passing process through the sparse relational structures. Applying sparsity regularization during training reduces minADE20 by 5.3%, 9.1%, 3.6%, and 5.7% on the ETH-UCY (overall), ETH-UCY (difficult), SDD, and NBA datasets, respectively.
Hyperedge numbers
@@
In Table IV, we compare the model performance with different hyperedge (i.e., group) numbers on the NBA dataset. It shows that setting 5 as the maximum number of hyperedges achieves the best performance based on a hyperparameter search. The model may not capture sufficient group-wise relations with only 2 hyperedges. At the same time, too many hyperedges (e.g., 11 and 110) may lead to overly dense group-wise relations that cannot capture clear group behaviors. In Table I-III, the reported numbers of SHG show the best performance after doing a hyperparameter search. Based on the experimental results, we find that using a very large number of hyperedges in SHG leads to serious overfitting. Moreover, the standard deviation of the prediction error increases as the maximum hyperedge number increases because a larger number of hyperedges may lead to more local optima, which enlarges the randomness of the model performance. These observations imply that although pairwise relations are a special type of group-wise relations where only two agents are involved, it is not sufficient to just use a large number of hyperedges to cover all the relations, which motivates our design of the integrated pairwise and group-wise relational reasoning.
Edge and hyperedge types
@@
In Table V, we compare the model performance with different numbers of edge and hyperedge types (i.e., relation types) on the NBA dataset. It shows that setting 3 as the edge and hyperedge type number achieves the best performance based on a hyperparameter search. The model may not capture a sufficient variety of pairwise or group-wise relations with only 2 edge and hyperedge types while too many types (e.g., 6) may lead to overfitting. We can see in Table V that the standard deviation of the prediction error increases as the type number increases because a larger variety of relations may lead to more local optima, which enlarges the randomness of the model performance.
Attention mechanisms
@@
To validate the effectiveness of the attention mechanisms in Eq. (5) and (10), we replace the attention weights in DCG+DHG+SM+SH+SP with equal weights, which shows that adopting the attention mechanism reduces minADE20 by 14.3%, 16.7%, 13.8%, and 11.5% on the ETH-UCY (overall), ETH-UCY (difficult), SDD, and NBA datasets, respectively. This implies the benefits of modeling different contributions of agents to group-wise relations and the influence of different groups on a certain agent.
@@ VI-E3 Qualitative Analysis
Method | SR | CR | TR | NT | PL | ITR | SD | GI |
ORCA [67] | 80.771.93 | 18.591.82 | 0.640.10 | 14.980.09 | 20.710.36 | 18.280.82 | 0.360.02 | 22.451.51 |
DSOA[68] | 90.250.98 | 4.200.68 | 5.550.32 | 16.020.11 | 23.150.42 | 27.411.01 | 0.180.02 | 38.412.12 |
RL (no prediction) | 83.022.41 | 16.632.33 | 0.350.08 | 16.390.16 | 22.560.21 | 7.210.27 | 0.320.03 | 9.910.20 |
RL (constant velocity) | 84.112.55 | 15.631.31 | 0.260.05 | 15.860.18 | 21.710.20 | 4.010.11 | 0.290.02 | 8.980.12 |
Group-Aware [42] | 91.681.54 | 7.931.45 | 0.390.06 | 15.270.11 | 21.080.19 | 3.420.23 | 0.330.03 | 8.790.21 |
CrowdNav++ [13] | 93.211.31 | 6.661.30 | 0.130.03 | 14.810.07 | 20.370.15 | 3.110.10 | 0.380.02 | 8.660.14 |
Ours (w/o DHG+GR) | 94.571.21 | 5.331.33 | 0.100.03 | 14.880.08 | 20.390.14 | 3.120.10 | 0.360.03 | 8.510.18 |
Ours (w/o GR) | 96.080.61 | 3.810.55 | 0.110.03 | 14.780.10 | 20.330.15 | 3.090.10 | 0.370.02 | 8.430.12 |
Ours | 96.310.53 | 3.590.50 | 0.100.03 | 14.920.15 | 20.530.13 | 3.080.08 | 0.400.02 | 5.210.11 |
Oracle | 97.280.46 | 3.070.41 | 0.050.05 | 14.570.24 | 20.050.13 | 3.010.09 | 0.420.03 | 4.290.04 |
@@ VI-F1 Quantitative Analysis
By employing our trajectory prediction model with relational reasoning, the social navigation method significantly outperforms the rule-based method ORCA [67], the dynamical system-based obstacle avoidance (DSOA) algorithm [68], the non-predictive method RL (no prediction), the predictive baseline RL (constant velocity), Group-Aware [42], as well as CrowdNav++ [13] across all the evaluation metrics. The improvement is evident in achieving a higher success rate (SR), lower collision rate (CR) and timeout rate (TR), and shorter navigation time (NT) and path lengths (PL). Meanwhile, our method notably enlarges the social distance from humans and reduces improper intrusions into human personal or group spaces, as indicated by SD, ITR, and GI. This enables the robot to find safer navigation trajectories and minimize interference with pedestrians. Although the DSOA algorithm [68] effectively prevents collisions between robots and individual humans or small numbers of pedestrians, its disregard for group dynamics may cause the robot to inadvertently enter the center of a group. This positioning can lead to longer navigation paths as the robot must then navigate out of the group, potentially resulting in unavoidable collisions and unsafe navigation conditions. Moreover, the Group-Aware model [42], while considering group formation of pedestrians, exhibits limited performance when interacting with multiple surrounding dynamically evolving groups, which may lead to suboptimal actions. Our method achieves comparable performance to the Oracle method which has access to the ground truth human trajectories, showing the effectiveness of our prediction model.
@@ VI-F2 Ablative Analysis
Group-wise reasoning
Ours (w/o DHG+GR) in Table VI represents a variant of our method with only pairwise reasoning and hence it disables group-based reward. Due to the lack of understanding of group relations, this variant leads to a higher CR and lower SD because the prediction model may not predict group behaviors accurately, which degrades the safety of robot navigation. Meanwhile, it causes a higher ITR and GI during navigation, which may make the surrounding humans less comfortable. Although it achieves a slightly shorter NT and PL (i.e., higher efficiency), the improvement is brought by the cost of social appropriateness, which is not desired.
Group reward
Ours (w/o GR) in Table VI represents a variant of our method without applying the group-based reward for RL policy learning. It still uses group-wise reasoning for trajectory prediction. The results show that without explicitly considering group relations in the reward function design, the robot policy still leads to a higher CR, ITR, GI, and a lower SD, although it achieves a comparable SR to our complete method. The NT and PL of our method are slightly larger because the policy tends to choose a safer and socially compliant trajectory which may take detours during navigation, which is acceptable and even desired in practice.
@@ VI-F3 Qualitative Analysis
VI-F4 Real-World Experiments
We deploy our model, trained in the simulation environments, directly on a RosMaster X3 robot equipped with four Mecanum wheels, which enables independent speed control in both horizontal and lateral directions. Different from ideal holonomic kinematics simulations, the RosMaster X3 robot cannot handle sudden changes in velocity. Therefore, we implement smoothness constraints to limit the maximum change in velocity, setting the maximum acceleration to and the maximum velocity to . We apply direct clipping after receiving the velocities generated by our RL agent. We use a 2D LiDAR, RPLIDAR-A1, to detect human positions through a learning-based human detector, DR-SPAAM [69]. The robot navigates on a large outdoor terrace with natural pedestrians, which covers an area larger than . In our environment, pedestrians walk at a natural speed with no constraints as shown in Fig. 9.
We select goal points for the robot close to human crowds to encourage maximum interactions. In our experiments, the robot is surrounded by up to 20 pedestrians in complex, dynamic scenarios. As shown in our experiment videos, and the robot successfully navigates to its goal while making appropriate decisions without collisions. In addition, we intentionally position humans to block the robot’s path, and the robot responds promptly to avoid collisions, indicating that our navigation policy adapts well to various pedestrian motion patterns and interacts with real humans proactively despite noises from the perception module and robot dynamics.
@@ VII Conclusion
In this paper, we present a systematic dynamic relational reasoning framework that infers pairwise and group-wise relations, which proves to be effective for multi-agent trajectory prediction and social robot navigation. In addition to inferring the conventional graphs that model pairwise relations between interacting agents, we propose to infer relation hypergraphs to model group-wise relations that exist widely in real-world scenarios such as human crowds. Different hyperedge types represent distinct group-wise relations. We employ a dynamic reasoning mechanism to capture the evolution of underlying relational structures. We propose effective mechanisms to regularize the smoothness of relation evolution, the sharpness of the learned relations, and the sparsity of the learned graphs and hypergraphs, which not only enhances training stability but also reduces prediction error. The proposed approach infers reasonable relations and achieves state-of-the-art performance on both synthetic crowd simulations and multiple benchmark datasets. Moreover, we present a deep reinforcement learning framework for social robot navigation that incorporates relational reasoning and trajectory prediction in a systematic way. It achieves state-of-the-art performance in terms of safety, efficiency, and social appropriateness compared with advanced baselines. A limitation of this work is that although our method outperforms the strongest baseline by a large margin, it still cannot achieve a zero collision rate in very dense scenarios. Therefore, in future work, we will investigate how to enable a provably safety guarantee.